OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

\[the equation is "Y-3X=-14"\] the steps: get the slope of the given line "m" ; then the slope of the needed equation ="m" also ;then Y-Y1=m(X-X1) ;then you get the equation